OpenStudy (anonymous):

Math Analysis: Use an appropiate identity to find the exact value of the expression. 12. sin15 13. tan23 + tan22/1-tan23tan22

OpenStudy (anonymous):

12.) sin(45 - 30) = sin 45cos 30 - cos 45sin 30 = (√2/2)(√3/2) - (√2/2)(1/2) = (√6/4) - (√2/4) = (√6 - √2)/4

OpenStudy (campbell_st):

this uses sum and difference formula sin 15 = sin(45 - 30) = sin45cos30 - cos45 sin30 45 and 30 are exact values which can be substituted into the equations \[\sin45 = \cos(45) = \frac {1}{\sqrt{2}}\] \[\sin(30) =1/2 ===========\cos(30) = \frac{\sqrt{3}}{2}\]

OpenStudy (campbell_st):

the next question is the expansion of tan(23 + 22) = tan (45) = 1

OpenStudy (anonymous):

yeah, i got it thx :) nxt question!!! lol

OpenStudy (anonymous):

Math Analysis: Find the indicated trigonometric function for 0 given 0 is an angle in standard position with terminal side defined by the given point. (-5,12); find cos0

OpenStudy (anonymous):

wait i mean

OpenStudy (anonymous):

sin^20 + sin0 -1 cos^20